OpenStudy (anonymous):

giving more medals Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

jigglypuff314 (jigglypuff314):

volume of cube = (side)^3 = (side)*(side)*(side)